### v1

# Searching for quantum-gravity footprint around stellar-mass black holes

Abstract
According to the asymptotically safe gravity, black holes can have characteristics different from those described according to general relativity. Particularly, they are more compact, with a smaller event horizon, which in turn affects the other quantities dependent on it, like the photon ring and the size of the innermost stable circular orbit. We decided to test the latter by searching in the literature for observational measurements of the emission from accretion disk around stellar-mass black holes. All published values of the radius of the inner accretion disk were made homogeneous by taking into account the most recent and more reliable values of mass, spin, viewing angle, and distance from the Earth. We do not find any significant deviation from the expectations of general relativity. Some doubtful cases can be easily understood as due to specific states of the object during the observation or instrumental biases.
1 Introduction
The description of all physical phenomena in the Universe in one unified theory is the ultimate dream of theoretical physics. Despite of decades of theoretical research [1, 2], little has been obtained on the observational/experimental side. Any experimental or observational evidence of quantum gravity remains elusive still today [3]. Today more than ever, it is necessary to find some observational constrain, to try stopping the uncontrolled proliferation of theories.
Among the different available variants of quantum gravity theories, asymptotically safe gravity (ASG) could offer some interesting opportunities. First developed by Steven Weinberg [4], it has its pillars on quantum field theory and renormalization group [5, 6]. Most important, it is expected to have observational consequences on astrophysical and cosmological scales [7], and in particular around black holes [8, 9, 10, 11]. For large values of the distance $r>>l_{\rm Planck}$, where $l_{\rm Planck}\sim 1.62\times 10^{-33}$ cm is the Planck length, the gravitational constant $G(r)$ becomes [9]:
$G(r)\sim G_{\rm N}\left(1-\frac{\xi}{r^{2}}\right)$ (1.1)
where $G_{\rm N}\sim 6.67\times 10^{-8}$ cm3 g-1 s-2 is the classical gravitational constant, and $10<\xi<$ 1 is the coupling coefficient.
The main effect on black holes is a more compact object, although without curvature singularities [10]. A smaller event horizon affects also all the other related quantities such as the photon ring [10], the ergosphere [9] and the radius of the innermost stable circular orbit $r_{\rm isco}$ [11]. The effects of ASG should be larger for smaller masses, although the constraint of $10<\xi<$ 1 sets severe limits on them. When the mass of the compact object is much larger than the Planck mass – which is any real case – the critical surfaces of an ASG black hole should not be distinguishable from the classical one according to [8] (see their Fig. 4 and 5). However, [11] suggested that differences up $\sim 30$% in the $r_{\rm isco}$ are expected in the case of maximally rotating prograde black holes already with a small value of $\xi$ (see his Table I). Moreover, the constraint on $\xi$ was put to question in Eichhorn and Held [10], who claim that values greater than 1, which lead to larger deviations from GR, can be admitted within the framework of the ASG. The same authors also stress that “irrespective of theoretical considerations, any observational avenue to put constraints on deviations from GR, should be explored” [10]. This is also our philosophical approach to the problem, searching for any deviation from GR around black holes.
Some attempts have been already made based on X-ray spectroscopy [12, 13, 14], but to date no deviations were found. We focus our study on the search for deviations from the value of the $r_{\rm isco}$ in the case of stellar-mass black holes calculated according to general relativity. Since these objects are the smallest black holes (primordial black holes were never observed to date), the effects of ASG – if any – should be maximized. Anyway, it should be noted that the theoretical minimum mass of a black hole is estimated to be $\sim 2.33M_{\odot}$ [15], which is still $\sim 38$ orders of magnitude greater than the Planck mass (but remind the above quote by Eichhorn and Held [10]). In addition, since the peak of the emission from the accretion disk is inversely proportional to the square root of the mass of the compact object, it falls in the X-ray energy band, where are less problems of contamination from the nearby environment. In the case of supermassive black holes, the disk emission peaks in the ultraviolet, where there are contaminations from emission lines of the broad-line region and the host galaxy.
As a first step, since we need of observations of black holes in specific states (high/soft, see Sect. 3), we searched in all the available literature for measurements of the inner radius of the accretion disk. Then, we made all the values homogeneous to compare each other.
2 A short summary of the accretion disk theory and models
According to the standard theory [16, 17], the accretion disk around black holes is optically thick and geometrically thin, although there are other models, particularly in the case of very high or very low mass accretion rate [18]. From the observational point of view, the disk can be considered as a series of rings of increasing size going outward, each one emitting black body radiation with increasing temperature $T(r)$ as the radial distance $r$ from the black hole decreases (e.g. see Fig. 2 in [19]) according to:
$T(r)\propto r^{-\frac{3}{4}}$ (2.1)
This model is named multicolor accretion disk [20, 21, 22], and is implemented as diskbb model in the xspec software package [^2] https://heasarc.gsfc.nasa.gov/xanadu/xspec/. The highest temperature $T_{\rm in}$ corresponds to the innermost part of the disk, the one closest to the event horizon.
The disk luminosity $L_{\rm d}$ [erg s-1] can be calculated via the Stefan-Boltzmann law:
$L_{\rm d}=4\pi r_{\rm in}^{2}\sigma T_{\rm in}^{4}$ (2.2)
where $\sigma=5.67\times 10^{-5}$ erg cm-2 s-1 K-4 is the Stefan-Boltzmann constant, and $T_{\rm in}$ [K] is the innermost temperature at the radius $r_{\rm in}$ [cm]. This equation is valid when the outer radius of the accretion disk $r_{\rm out}$ is much larger than $r_{\rm in}$, so that the outer temperature $T_{\rm out}$ can be considered negligible.
It is worth noting that X-ray spectra cannot measure the innermost effective temperature, because of Comptonization. The observed color temperature $T_{\rm col}$ is greater than the effective one by a factor $\kappa\sim 1.7-2.0$, which depends on the accretion rate [23]. This affects also the effective location of the inner radius, which should be corrected by a factor $\varsigma=0.412$ [24]. Therefore, to use the observed quantities, Eq. (2.2) is rearranged as:
$L_{\rm d}=4\pi\sigma\left(\frac{r_{\rm in}}{\varsigma}\right)^{2}\left(\frac{T_{\rm col}}{\kappa}\right)^{4}$ (2.3)
The luminosity can be calculated by means of geometrical considerations, since the thin disk can be assumed to be flat:
$L_{\rm d}=2\pi\frac{d^{2}}{\cos i}F_{\rm bol}$ (2.4)
where $F_{\rm bol}$ is the disk bolometric flux [erg cm-2 s-1], $d$ is the luminosity distance [cm], and $i$ is the viewing angle of the disk [deg]. By combining Eqs. (2.3) and (2.4), it is then possible to estimate the inner radius of the accretion disk:
$r_{\rm in}=\varsigma\kappa^{2}d\sqrt{\frac{F_{\rm bol}}{2\cos i\sigma T_{\rm col}^{4}}}$ (2.5)
The normalization $N$ of the diskbb model – that can be estimated via spectral fitting – includes some of the above factors, and is equal to:
$N=\left(\frac{r_{\rm col}}{d_{10}}\right)^{2}\cos i$ (2.6)
where $d_{10}$ is the luminosity distance in units of $10$ kpc, and $r_{\rm col}=r_{\rm in}/\varsigma$ is the color radius, which must be corrected according to the above cited factors as:
$r_{\rm in}=\varsigma\kappa^{2}r_{\rm col}=\varsigma\kappa^{2}d_{10}\sqrt{\frac{N}{\cos i}}$ (2.7)
Despite its simplicity, the diskbb model performs reasonably well, as it results from the comparison with more detailed models [25, 19, 26, 27], although with some caveats [28, 19].
3 Gathering data
Eqs. (2.5) and (2.7) offer two possibilities to estimate the innermost radius of the accretion disk. In addition to the quantities measured via the fit of the X-ray spectrum ($N$, $T_{\rm col}$, $F_{\rm bol}$), it is necessary to know the distance $d$, the viewing angle $i$, but also the mass of the black hole $M$ and its spin $a$ to estimate the radius of the innermost stable circular orbit $r_{\rm isco}$ according to GR (Kerr metric, [17]):
$r_{\rm isco}=r_{\rm g}\left[3+Z_{2}-\sqrt{(3-Z_{2})(3+Z_{1}+2Z_{2})}\right]$ (3.1)
[…]
The values of $r_{\rm g}$, $a$, $d$, and $i$ changed often during time, depending on the instrument evolution and the observational constrains (e.g. finding the stellar companion is often difficult, given the presence of many molecular clouds and interstellar dust). We adopted the latest values, which are not always consistent with those adopted by many other authors when converting their observational data. For example, many authors did not report the normalization $N$ of the diskbb model, but only the calculated $r_{\rm in}$ or even $r_{\rm col}$. Therefore, to extract a value comparable with other observations made at different epochs and with different instruments, it is necessary to know which values of $d$ and $i$ were adopted in the conversion. This is not granted: sometimes the values used in the conversion were not written in the paper, which made it impossible to recalculate the original normalization. Other problems encountered when searching in the literature were: missing measurement errors, missing units of measurement, plain errors and/or typos. In case of errors, it was often possible to correct them by recalculating the described model. For example, when finding a weird normalization $N$, it is possible to recalculate it if $T_{\rm col}$ and $F_{\rm bol}$ were reported. Fluxes were often reported without measurement errors: in this case, a typical $10$% error was adopted.
Yet another issue regards the instruments. In the case of stellar-mass black holes, the color temperature of the multicolor accretion disk peaks in the range $\sim 0.1-1$ keV. However, many detectors have a low-energy threshold of a few keV, which made it difficult to correctly extrapolate to lower energies. For example, one of the most used and prolific satellites in the observation of X-ray binaries was the Rossi X-ray Timing Explorer (RXTE) [^3] https://heasarc.gsfc.nasa.gov/docs/xte/xtegof.html. The Proportional Counter Array (PCA) worked in the $2-60$ keV energy range, and its low-energy threshold degraded up to $3$ keV with the age. The $2$ keV threshold makes it difficult to correctly extrapolate a $T_{\rm col}=0.4$ keV, because only a small part of the spectrum can be detected, as shown in Fig. 1.
Figure 1: Unfolded spectrum of the diskbb model with $T_{\rm col}=0.4$ keV, and assuming a typical Galactic hydrogen column $N_{\rm H}=10^{21}$ cm-2. The vertical dashed line represents the $2$ keV low-energy threshold of RXTE/PCA. The flux is in arbitrary units (diskbb normalization set to 1).
In addition to the above cited problems, it is necessary to take into account that the accretion disk does not extend always down to the innermost stable circular orbit. Stellar-mass black holes change their activity from high/soft to low/hard states, and vice versa, passing through some intermediate states, depending on the accretion rate [37, 38, 39, 40]. High/soft states occur at high accretion rates. The X-ray spectrum is dominated by the thermal emission of a geometrically-thin optically-thick accretion disk, which can extend down to $r_{\rm isco}$, and it accounts for more than $50$% of the bolometric flux [38]. At low accretion rates (low/hard states), the thin disk is weak and cool, and a mildly relativistic radio jet is present. The X-ray spectrum is dominated by the optically-thin emission from the corona, modelled with a hard power law or a more detailed Comptonization model. The intermediate state is an obvious mixture of the two above cited states. There are also many other variations, distinguished by small details [^4] See the plot: https://www.issibern.ch/teams/proaccretion/Documents.html. However, for our purposes, we focus our attention on high/soft states, when the accretion disk extends downward the innermost stable circular orbit. This is also a requirement to have a reliable measurement of $r_{\rm in}$, because the multicolor disk model tends to underestimate the inner radius when the corona significantly contributes to the bolometric flux [28], as it happens in the intermediate states, for example.
4 Results
4.1 Cygnus X-1
This is the oldest and best known stellar-mass black hole. It has the most reliable measured values of $M=21.1_{-2.3}^{+2.2}M_{\odot}$, $d=2.22_{-0.17}^{+0.18}$ kpc, $i=27.51_{-0.57}^{+0.77}$ deg, and $a=0.9696-0.9985$ [41]. It is then possible to calculate $r_{\rm g}=31\pm 3$ km and $r_{\rm isco}=(1.21-1.74)r_{\rm g}$. The innermost radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 2.
Figure 2: Cygnus X-1: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR. Reference number refers to the source of data: 1: [42]; 2: [43]; 3-4: [44]; 5: [45]; 6: [46]; 7: [47]; 8-11: [48]; 12-13 [49]; 14: [50].
The smallest value of $r_{\rm in}=(1.29_{-0.27}^{+0.28})r_{\rm g}$ – still consistent with the expectations from GR – was derived from a simultaneous observation with Suzaku and NuSTAR ($1-300$ keV) between October 31 and November 1, 2012 [46]. We considered the best fit model 4 in Table 2, where the normalization of diskbb is reported to be $N=20800_{-800}^{+1200}$, from which we can estimate $r_{\rm in}$ via Eq. (2.7). Unless explicitly stated otherwise, we always assumed $\kappa=1.7$.
4.2 GRS 1915+105
This is one of the most monitored and studied stellar-mass black hole. Also in this case, there are reliable measured values of $M=11.8\pm 0.6M_{\odot}$, $d=9.4\pm 1.0$ kpc, $i=64\pm 4$ deg, and $a=0.970-0.997$ [51, 52, 53]. It is then possible to calculate $r_{\rm g}=17.4\pm 0.9$ km and $r_{\rm isco}=(1.28-1.74)r_{\rm g}$. The innermost radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 3.
Figure 3: GRS 1915+105: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR. Reference number refers to the source of data: 1-2: [54]; 3: [55]; 4-7: [56]; 8-10: [57]; 11: [58]; 12: [59]; 13-22: [60]; 23-27: [61]; 28-29: [62]; 30-32: [63]; 33-36: [64]; 37-46: [65]; 47-50: [66]; 51-54: [67]; 55-60: [68]; 61: [69]; 62: [70]; 63-68: [71]; 69: [72].
This is the only case where we adopted $\kappa=1.9$, as suggested by [65, 69], since the color temperature is generally higher than usual for a stellar-mass black hole ($T_{\rm col}\sim 1-2$ keV).
Also in this case, $r_{\rm in}$ is consistent with the expectations from general relativity within the measurement errors. Some cases are borderline: one reason can be that the diskbb does not include relativistic corrections (gravitational redshift and Doppler boosting), which can become significant for high inclinations [19]. We remind that $i=64\pm 4$ deg for GRS 1915+105 [53].
4.3 XTE J1550-564
This black hole is interesting, because it is not maximally rotating. Its spin has been estimated to be $a=0.49_{-0.20}^{+0.13}$ [73]. The other quantities necessary to estimate the reference and measured values of the inner radii are $M=9.10\pm 0.61M_{\odot}$, $d=4.38_{-0.41}^{+0.58}$ kpc, and $i=74.7\pm 3.8$ deg [74]. The gravitational radius is $r_{\rm g}=13.4\pm 0.9$ km, and $r_{\rm isco}=(3.74-5.01)r_{\rm g}$. The innermost radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 4.
Figure 4: XTE J1550-564: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR. Reference number refers to the source of data: 1: [75]; 2-13: [76]; 14-15: [77]; 16: [78]; 17: [79]; 18-20: [80]; 21: [81]; 22: [82]; 23-27: [83].
In this case, there are more significant deviations (reference numbers 14, 15, 18, 21). The cases 14 and 15 were from [77] and refer to RXTE observations in the $3-200$ keV energy band between April and June 2000. Table 1 of the cited paper reported the apparent radius calculated from the normalization of diskbb by assuming $d=6$ kpc and $i=73.1$ deg. The fits were done every one-two days and a thermal component is reported in 21 over 43 fits, with $T_{\rm col}\sim 0.37-0.95$ keV. We can already note that such temperature can be biased by the $3$ keV low-energy threshold of the detector. Since we are interested in searching for deviations from the expectations of general relativity, we selected the two smallest apparent radii, which are $r_{\rm col}\sim 32.5$ km (MJD $51662.2$), and $r_{\rm col}\sim 39.4$ km (MJD $51673.4$). There is also a third value of $r_{\rm col}\sim 39$ km at MJD $51682.3$, but the photon index of the power-law model is harder than those of the previous two cases ($1.76$ vs $2.33$, and $2.31$). This suggests the presence of a significant Comptonization, which can alter the estimate of the apparent radius [28], and therefore we did not consider it. For the two remaining cases, it is necessary to understand if they were measured during a high/soft state, with negligible corona contribution. As already written, the photon indexes of the power-law models were steep, but there are no indication of the fluxes due to the different components. We found some indications only in Fig. 4 in [77], which plots the fluxes in the $2-50$ keV energy band of the two models (power law vs diskbb). One point corresponding to MJD $51662.2$ is explicitly indicated: in this case, the power-law flux is about $2-3$ times that from the accretion disk. The other point, corresponding to MJD $51673.4$, is not indicated. However, by looking at the other points in the plot, one can reasonably think that the power-law component should have been at least comparable to the thermal emission. Therefore, we can conclude that these two inner radii estimates do not refer to high/soft states, and are affected by a strong Comptonization, which led to an underestimation of the value [28]. Similar cases are the number 18 from [80] (it is the period 7 in Table 1), and the number 21 from [81] (it is the case C, model 2 in Table 2).
A last note of caveat: recently, the viewing angle has been challenged by [82]. They suggested a value of $i\sim 40$ deg, much smaller than previously measured. This would have a significant impact on the estimation of the inner radius. The value calculated for $i\sim 75$ deg is $r_{\rm in}\sim 5.90r_{\rm g}$ (reference number 22), but in the case of $i\sim 40$ deg, then $r_{\rm in}\sim 3.46r_{\rm g}$, smaller than expected from general relativity ($r_{\rm isco,min}\sim 3.74r_{\rm g}$). This new viewing angle seems quite unlikely, because the kinematic analysis of the relativistic jet sets a tight constrain on the viewing angle, indicating a value of $i\sim 71^{\circ}$ [84].
Figure 5: GX 339-4: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR. Reference number refers to the source of data: 1-2: [86]; 3: [87]; 4-5: [88]; 6: [89]; 7: [90]; 8-10: [91]; 11-13: [92]; 14-18: [93]; 19: [94]; 20: [95]; 21: [96]; 22: [97]; 23: [98]; 24-25: [99]; 26: [100]; 27-29: [101]; 30-33: [102]; 34-36: [103]; 38-40: [105]; 41: [106]; 42: [107]; 43: [108].
4.4 GX 339-4
Yet another well-known and observed stellar-mass black hole. Its reference quantities are: $M=9.0_{-1.2}^{+1.6}M_{\odot}$, $d=8.4\pm 0.9$ kpc, $i=30\pm 1$ deg, and $a=0.95_{-0.08}^{+0.02}$ [85]. The gravitational radius is $r_{\rm g}=13\pm 2$ km, and $r_{\rm isco}=(1.73-2.51)r_{\rm g}$. The innermost radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 5.
Also in this case, there are no significant deviations. The only value outside the expected range – but stil consistent within the measurement errors – is the number 14 [93]. It refers to a INTEGRAL and XMM-Newton observation ($0.7-200$ keV) performed between January and March 2007. The value of $r_{\rm in}=1.21_{-0.29}^{+0.98}r_{\rm g}$ has been estimated from the data of Epoch 1 of Table 2. We noted the presence also of a hard component modelled with a power law with $\Gamma\sim 1.46$, and Table 3 stated that the disk component accounts for $\sim 21$% of the total flux. Again, a significant Comptonization led to an underestimate of the inner radius of the disk.
4.5 XTE J1650-500
As we wrote in the Introduction, we expect that the most significant deviations from GR may occur around the smallest black holes. Therefore, we decided to study also two objects with the smallest mass, although with large uncertainties and the lack of some reference information. The first one is XTE J165050016505001650-5001650 - 500 and has its reference quantities are: $M=4.0\pm 0.6M_{\odot}$, $d=2.6\pm 0.7$ kpc, $i=70\pm 4$ deg, and $a\sim 0.9982$ [109, 110, 111]. The gravitational radius is $r_{\rm g}=5.9\pm 0.9$ km, and $r_{\rm isco}\sim 1.23r_{\rm g}$.
We found only two estimates in the literature, both showing no deviations from the expected value of $r_{\rm isco}$: $r_{\rm in}=(18\pm 8)r_{\rm g}$ [112], and $r_{\rm in}=(5.3\pm 1.7)r_{\rm g}$ [113].
Searching for some unpublished data in the archives, we found one Swfit observation (id 00031762001) performed on July 17, 2010. We analysed the data by using HEASoft v. 6.33.2 software package and CALDB updated on August 28, 2024. Swift/XRT data were reduced by using standard procedures (xrtpipeline with default values). We found no sources, with a flux upper limit of $\sim 3\times 10^{-11}$ erg cm-2 s-1 ($3\sigma$ detection, $\sim 1.0$ ks exposure in photon counting mode, $N_{\rm H}=3.68\times 10^{21}$ cm-2 [114], power-law model with $\Gamma=2$). This means that XTE J1650-500 was likely in a quiescent state.
4.6 GRO J0422+32
The last object is the candidate to be the smallest black hole known, with $M=2.7_{-0.5}^{+0.7}M_{\odot}$, $d=2.49\pm 0.30$ kpc, $i=55.6\pm 4.1$ deg [115, 116]. No information about the spin was found, and therefore it is not possible to estimate the expected $r_{\rm isco}$. We found only one paper useful to estimate the inner radius: $r_{\rm in}=(5.1\pm 2.3)r_{\rm g}$ [117]. We also found one Swift observation (id 00032976001) performed on October 3, 2013. We analysed the data with the same procedure adopted for the above cited case, but found no source, with a a flux upper limit of $\sim 1.5\times 10^{-11}$ erg cm-2 s-1 ($3\sigma$ detection, $\sim 1.2$ ks exposure in photon counting mode, $N_{\rm H}=1.51\times 10^{21}$ cm-2 [114], power-law model with $\Gamma=2$). Also GRO J0422+32 was likely in a quiescent state.
5 Conclusions
We searched in all the available literature for observations of X-ray binaries in high/soft state, where it is expected that the inner disk is closest to the innermost stable circular orbit ($r_{\rm in}\sim r_{\rm isco}$), to test if there are deviations from general relativity. All the measurement were made homogeneous according to the most recent and reliable values of the reference quantities ($r_{\rm g}$, $a$, $d$, $i$). All the measured radii are consistent with the expectations of general relativity. A few anomalous cases can easily be reconciled by taking into account the Comptonization and how it affects the estimation of $T_{\rm in}$, the uncertainties on the reference quantities, and the instrumental biases.
[…]
$\tilde{\xi}=\frac{\xi}{r_{\rm g}^{2}}\gtrsim 0.028$ (5.1)
We note that a change in $G$ would affect also the measurement of the mass of the black hole. In the case of XRBs, the mass is measured via the Kepler’s third law, which includes $G$ [118]. Therefore, we should revise the mass of Cyg X-1 by using Eq. (1.1) in the mass function equation (cf Eq. 1 in [118]). However, the orbit of the binary system (the semimajor axis is $\sim 0.244\,\mathrm{AU}\sim 3.7\times 10^{7}$ km, [41]) is much greater than the size of $r_{\rm isco}=1.61r_{\rm g}\sim 50$ km, and, given a dependence on $r^{2}$ of $G(r)$ (cf Eq. 1.1), the effect is negligible and smaller than the present measurement error.
We cannot set any constrain on negative values of $\xi$ with this method. In this case, $r_{\rm isco}$ from ASG should be greater than the value from GR. However, a greater $r_{\rm in}$ is expected from state transitions in stellar-mass black holes. Particularly, in the hard state, the inner disk is truncated and $r_{\rm in}$ is of the order of tens $r_{\rm g}$ (see, however, [119]).
Other, more complex expressions of $G$ are currently under study [120], and will be the topic of another essay.
Before concluding, we would like to underline the need to improve the measurements of the reference quantities, particularly the spin, which are crucial to calculate the GR values to be compared with. It is also necessary to better understand how the Comptonization affects $T_{\rm in}$. Last, but not least, more observations on the smallest black holes are needed.
Acknowledgments
LF thanks Roberto Della Ceca, director of the Brera Astronomical Observatory – INAF, for partially funding this research, and Daniele Malafarina for useful questions after the presentation of this work at the IV International FLAG workshop: the Quantum and the Gravity, (Catania, Italy, 9-11 September 2024).